OpenStudy (anonymous):

Find the smallest number that has the factors of 2, 3,4, 5, 6, 7, 8, 9, and 10

Parth (parthkohli):

That's a fancy way to say lowest common multiple. Do you know how to calculate the lowest common multiple?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

doi multiply all of them #s

OpenStudy (anonymous):

the*

Parth (parthkohli):

Oh, well... suppose that you're given the list of numbers. The lowest common multiple of all those numbers is the least numbers that is a multiple of all those numbers in the list (i.e., occurs in the multiplication table). Multiplying all those numbers is a good way to ensure that the number obtained will be a common multiple, but it is not necessary that the number will be the lowest common multiple.

Parth (parthkohli):

I'm too lazy to explain the method used to obtain the lcm, so I'll just link you.

OpenStudy (anonymous):

well add all the numbers and devide it by how many numbers are listed.

Parth (parthkohli):

what

OpenStudy (anonymous):

divide sorry

OpenStudy (anonymous):

not devide....

OpenStudy (anonymous):

add them??

Parth (parthkohli):

Suppose that you want to find the lcm of 9, 8, 10. Step 1: List out the prime factorisation of all the numbers, using exponents when necessary. For example, \(10 = 2\cdot 5\), \(9 = 3^2\), \(8 = 2^3\). Step 2: List out all the unique factors you see. \(2, 5, 3\) Step 3: Raise the factors to the highest exponent you see in all the prime factorisations. \(2^3, 5^1, 3^2\) Step 4: Multiply all the numbers you obtain in step 3.

OpenStudy (rational):

`1 = 1` 2 = 2^1 3 = 3^1 4 = 2^2 `5 = 5^1` 6 = 2*3 `7 = 7^1` `8 = 2^3` `9 = 3^2` 10 = 2*5